OpenStudy (worldunseen):

If g(x) = x^(2/3), show that g'(0) does not exist.

OpenStudy (mathmate):

Have you calculated g'(x) yet?

OpenStudy (worldunseen):

@mathmate isn't that just \[g'(x) = \frac{ 2 }{ 3*\sqrt[3]{x} }\]

OpenStudy (worldunseen):

because of the power rule

OpenStudy (irishboy123):

\(\Large \checkmark\)

OpenStudy (worldunseen):

@IrishBoy123 Oh and then it g'(0) would be undefined.

OpenStudy (mathmate):

@WorldUnseen exactly, well done! :)

OpenStudy (worldunseen):

Wow thanks everyone! @mathmate @IrishBoy123

OpenStudy (irishboy123):

sorta the derivative is the limit and here you come at it from different sides and you get different answers so it's not that simple

OpenStudy (worldunseen):

@IrishBoy123 Yeah, I know, that's how it really should be done. I just don't know how to take the limit of that on both sides.

OpenStudy (irishboy123):

plotting it is always a good idea

OpenStudy (worldunseen):

@IrishBoy123 I know what the graph looks like, and it's pretty obvious why it can't exist, I just wanted to see how to explain it with limits/derivatives

OpenStudy (irishboy123):

well, its a 2 sided limit different as you approach from either side so the limit does not exist

OpenStudy (worldunseen):

@IrishBoy123 Could you explain to me how to get the exact values of those limits?

OpenStudy (worldunseen):

I know it should have to be as h approaches 0 on both sides (+ and -), but I don't really know how to simplify from there.

OpenStudy (irishboy123):

\(\Large \lim\limits_{x \to 0^-} \frac{ 2 }{ 3*\sqrt[3]{x} } = -\infty \) \(\Large \lim\limits_{x \to 0^+} \frac{ 2 }{ 3*\sqrt[3]{x} } = \infty \)

OpenStudy (worldunseen):

@IrishBoy123 okay I see, thanks

OpenStudy (irishboy123):